# Click Without Compromise: Online Advertising Measurement via Per User Differential Privacy

### v1

@@ 3. Related Work
User level privacy guarantee is required. Two definitions of differential privacy exist in the advertising setting: event-level and user-level. The former safeguards individual events, while the latter conceals all events associated with a particular user across the entire ads measurement campaign. For instance, event-level privacy protects a single user’s click or purchase activity (i.e., conversion), whereas user-level privacy shields all user clicks or conversions. The chosen privacy level impacts the degree of perturbation applied. In this paper, we focus on studying the challenging and more stringent setting of user privacy. Choquette-Choo et al. introduced the matrix mechanism as a solution for user-level differential privacy in their work (choquette2022multi, ; choquette2024amplified, ). Their primary application is differentially private stochastic gradient descent. Nevertheless, it’s worth noting that these methods necessitate a minimum time gap of at least $b$ days between two records belonging to the same user. This requirement presents a challenge in the context of online advertising measurement systems, where a user can potentially have multiple conversions within a single day. Investigating user-level differential privacy in the context of continual observation, (dong2023continual, ; zhang2023differentially, ; george2022continual, ) used tree-based mechanisms to tackle the problem. Although effective for addressing the general streaming problem, these approaches do not leverage the specific structures inherent in advertising measurement systems. For example, in a general streaming process, a single user may contribute most of the records, leading to a large sensitivity (cropping the data related to this user will lead to a high bias). However, in an advertising system involving millions of users, it’s improbable for any individual user to make a substantial contribution to the dataset. In reality, most users generate only a small number of conversions. The removal of data associated with a few users is unlikely to have a significant impact on overall performance.
4. Differentially Private Measurement Design
We will begin by addressing the scenario involving a single publisher. In cases where multiple publishers are involved, we show how to extend the mechanism in Section 4.6. Given a streaming dataset ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ on a publisher for a specific advertiser, where ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[i]$ represents the number of conversions on day $i$, we are interested in releasing a differentially private prefix-sum. For example, given a streaming dataset ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}=({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[1],{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[2],{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[3],{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[4])$, we want to release $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[1],{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[1]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[2],{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[1]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[2]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[3],{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[1]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[2]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[3]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[4])$, which can be represented as a result of matrix-vector product $\mathbf{W}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$, where $\mathbf{W}=\left[\begin{smallmatrix}1&0&0&0\\ 1&1&0&0\\ 1&1&1&0\\ 1&1&1&1\end{smallmatrix}\right]$ is called the prefix-sum matrix.
Definition 4.1 (Prefix-sum Matrix).
For a given domain size $n$, a matrix $\mathbf{W}\in\mathbb{R}^{n\times n}$ is called prefix-sum matrix if
$\mathbf{W}[i,j]=\begin{cases}1,&1\leq i\leq j\leq n\\ 0,&\text{otherwise}\end{cases}.$
The next step is to design constraints specific to advertising measurement systems, aimed at preventing improbable scenarios from occurring. This targeted approach enhances accuracy by ensuring that the system operates within realistic bounds, reflecting the unique characteristics of advertising data.
Recalling from Section 2.2.1, the definition of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}^{\text{Global}}}}$ imposes a global limit on user contributions, overlooking the structures inherent in advertising measurement systems. Within these systems, users with particular interests typically exhibit consistent interactions with ads over time. For instance, sudden bursts of interest followed by complete disengagement are uncommon and may even signal malicious behavior. Instead, it is more typical for a user’s interest in a product to persist, resulting in conversions across multiple days. To capture this behavior effectively, we introduce the concept of bounded per-day contributions.

### v2

@@ 3 Related Work
User-level privacy guarantee is preferred. Two definitions of differential privacy exist in the advertising setting: event-level and user-level. The former safeguards individual events, while the latter conceals all events associated with a particular user across the entire ads measurement campaign. For instance, event-level privacy protects a single user’s click or purchase activity (i.e., conversion), whereas user-level privacy shields all user clicks or conversions. The chosen privacy level impacts the degree of perturbation applied. Choquette-Choo et al. introduced the matrix mechanism as a solution for user-level differential privacy in their work [14, 13]. Their primary application is differentially private stochastic gradient descent. Nevertheless,these methods necessitate a minimum time gap of at least $b$ days between two records belonging to the same user. This requirement presents a challenge in the context of online advertising measurement systems, where a user can potentially have multiple conversions within a single day. Investigating user-level differential privacy in the context of continual observation, [19, 56] used tree-based mechanisms to tackle the problem. Although effective for addressing the general streaming problem, these approaches do not leverage the specific structures inherent in advertising measurement systems. For example, in a general streaming process, a single user may contribute most of the records, leading to a large sensitivity (cropping the data related to this user will lead to a high bias). However, in an advertising system involving millions of users, most users generate only a small number of conversions in practice . The removal of data associated with a few users is unlikely to have a significant impact on the overall dataset. In this paper, we focus on the challenge of achieving the more stringent user-level differential privacy while satisfying the constraints and properties of an online ad measurement system.
4 Differentially Private Measurement Design
We design our system to satisfies the following properties. Flexibility. The goals vary across different types of ad campaigns, and we want to avoid redesigning the system for each specific case. Ideally, the system should accommodate various objectives with only minor modifications. Simplicity. The system should be easy to deploy and should introduce minimal changes to the existing non-private version. Adaptability. The system should automatically adjust to different datasets, ensuring the algorithm performs well across various types of data. Given these requirements, we propose the following mechanism design:
  1. 1.
    Each day $i$, we add random Gaussian noise $z_{i}$ to obtain the noisy result $\tilde{x}_{i}=x_{i}+z_{i}$, where $z_{i}\sim N(0,\sigma_{i}^{2})$ is a Gaussian random variable with mean zero and variance $\sigma_{i}^{2}$. While simple, adding independent noise—combined with per-day contribution limits—proves more effective in our experiments than more complex methods like the matrix mechanism.
  2. 2.
    Depending on the specific objective function, we initialize the scale parameters $\bar{\sigma}_{i}^{2}$, providing flexibility for different ad campaign goals.
  3. 3.
    Each day, we privately estimate the user contribution limit $r_{i}$. A user can contribute up to $r_{i}$ conversions on day $i$, with any additional conversions removed.
  4. 4.
    We then adjust the actual noise scale as $\sigma_{i}=r_{i}*\bar{\sigma}_{i}$, based on the contribution limit $r_{i}$ for day $i$. This data-dependent noise scaling makes the system more adaptive.
1:  Input: Target workload query matrix $\mathbf{Q}$; Objective function OBJ.
2:  Output: Noisy query answer.
3:  $\bar{r}_{1},\cdots,\bar{r}_{n}$ $=1$.
14:  $\bar{\sigma}_{1},\cdots,\bar{\sigma}_{n}$ = InitScales($\mathbf{Q}$, OBJ, $\{\bar{r}_{i}\}_{i=1}^{n}$).
25:  bound_list = [].
6:  for $i$ in range($n$do
7:     if $i<l$ then
8:        $r_{i}$= PrivateQuantile($D_{i}$, percent).
9:     else
10:        $r_{i}$ = UpdateBoundSVT($D_{i}$).
11:     end if
12:     Store the bound: bound_list.append($r_{i}$).
13:     Clip data: ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[i]$ = Clip($D_{i}$, $r_{i}$).
14:     Update scale: $\sigma_{i}=\bar{\sigma}_{i}/\bar{r}_{i}*r_{i}$.
15:     Add noise to data: $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}[i]={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[i]+N(0,\sigma_{i}^{2})$.
16:     Get query answer: $y_{j}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}^{T}\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}$.
317:  end for
18:  Return.
Algorithm 1 Ads Measurement with Bounded Per Day Contributions (Ads-BPC)
In the following sections, we describe each of the above components in detail. Algorithm 1 provides an overview of the AdsBPC mechanism. We begin by initializing the scale parameters based on the given query workload and objective function, with Section 4.3 covering the initialization process in detail. For the first $l$ days, we estimate the user contribution limit $r_{i}$ using a differentially private quantile algorithm, as explained in Section 4.4. For the remaining days, we apply the Sparse Vector Technique (SVT) to determine whether an adjustment to the user contribution limit is needed. Section 4.5 provides the implementation details for the SVT algorithm. We will begin by addressing the scenario involving a single publisher. In cases where multiple publishers are involved, we show how to extend the mechanism in Section 4.6. Before we introduce details of our algorithm, we show how to measure the utility of the proposed mechanism in Section 4.1, and give privacy properties of the mechanism in Section 4.2
4.1 Mechanism Utility
Recall that there are $n$ days in total, and we represent the data vector as $\mathbf{x}\in\mathbb{R}^{n}$, where $\mathbf{x}[i]$ represents the sum of attribution weights for a specific publisher on day $i$ ($1\leq i\leq n$). In order to analyze our mechanism using Theorem 2.6, we rewrite the noise addition step as $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}[i]={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[i]+N(0,\sigma_{i}^{2})=\sigma_{i}(\frac{1}{\sigma_{i}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[i]+N(0,1))$. That is, we add a Gaussian random variable with zero mean and variance of one to a scaled version of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[i]$. Let $\mathbf{B}=\text{diag}(\frac{1}{\sigma_{1}},\cdots,\frac{1}{\sigma_{n}})$, the private mechanism is defined as,
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}})=\mathbf{B}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+N(0,\mathbf{I}).$ (1)
We represent the $j$-th query vector as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}=({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}[1],{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}[2],\cdots,{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}[n])$, the true answer to this query is ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}^{T}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}$ = ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}[1]{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[1]+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}[2]{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[2]+\cdots+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}[n]{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[n]$. Because we perturb the ground truth data with independent Gaussian noise, $\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}[i]={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[i]+N(0,\sigma_{i}^{2})$, the noisy answer to the query ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}$ is ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}^{T}\tilde{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}}=\sum_{i=1}^{n}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}[i]({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[i]+N(0,\sigma^{2}_{i}))$ $=\sum_{i=1}^{n}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}[i]{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}[i]+\sum_{i=1}^{n}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}[i]N(0,\sigma^{2}_{i})$ $={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}^{T}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}+N(0,\sum_{i=1}^{n}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}[i]^{2}\sigma^{2}_{i}).$ We can see that the variance of the query ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}$ is
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{Var}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j};{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}):=\sum_{i=1}^{n}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j}[i]^{2}\sigma^{2}_{i}.$ (2)
The variance function of the mechanism ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}$ is defined as ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{Var}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{1};{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}),\dots,{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{Var}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{m};{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}}))$, here ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}}}$ is a function of ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{Var}}}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{q}}}_{j};{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}}})$. For example, ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}}}$ can be the sum function, then it represents the sum of variances for all queries.
4.2 Mechanism Privacy Budget
As stated in Theorem 2.6, the privacy budget is proportional to the $L_{2}$ sensitivity of the mechanism. Since we are adding independent noise to the data ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}}}_{i}$ each day, a natural way to reduce sensitivity is to limit the number of contributions a user can make daily. To achieve this, we introduce the concept of bounded per-day contributions.